The manager at GetFit notices that the suppliers have different levels of service. The levels of service of the companies are listed in Table 2a. The costs per % level of service depends on the SKU, and are listed in Table 2b. The costs and maximum total number of units that can be supplied by each supplier remain the same as in Part 1.

Please normalize the costs to Supplier A's level of service (92%) to get the answer.

Table 2a. Level of service (%)

Supplier A Supplier B Supplier C Supplier D
92 98 95 97

Table 2b. Cost per % level of service

Parts Cost per %
W32A 10
Y33A 2
W32B 1
X33B 3
Z32A 8

Considering the levels of service of the suppliers, what is the minimum total procurement cost for GetFit?
Answer #1

Using the reference of the previous subpart-

GetFit is a company that makes high-end exercise machines and the machine Pilates Power is their most popular product. For each production cycle of the Pilates Power machines, GetFit purchases 5 SKUs: W32A, Y33A, W32B, X33B, and Z32A. There are four suppliers: Supplier A, B, C, and D. The number of units needed in each production cycle and the prices for each unit offered by the suppliers are listed in Table 1a. The maximum total number of units that each company can supply are listed in Table 1b. For example, if Supplier C supplies 24800 pieces of SKU Z32A, it cannot supply anything else.

Table 1a. Number of units needed and unit prices (in dollars) offered by the suppliers.

Parts Number of units needed Supplier A Supplier B Supplier C Supplier D
W32A 9000 500 538 570 566
Y33A 16000 109 108 116 117
W32B 5000 27 28 30 29
X33B 31000 230 225 207 245
Z32A 33000 324 354 352 378

Table 1b. The maximum total number of parts that each company can supply

Supplier A Supplier B Supplier C Supplier D
25200 11200 24800 35800

Answer-

Let i =1,2,3,4,5 represent the parts W32A, Y33A, W32B, X33B, Z32A

Let j= A,B,C,D,E represent the supplier A, B,C,D,E

Let Pij be the integer variable representing the number of parts of part type i being supplied from j supplier

Let Cij be the cost of obtaining i part from j supplier

The Supplier B has 6% improvement, Supplier C has 3% improvement and Supplier D has 5% improvement in service level compared to Supplier A.

Therefore the part cost for any supplier = Cost of that part in Supplier A+(%improvement in Service Level*Cost per % level of service)

The Cost Matrix would thus be-

Supplier A Supplier B Supplier C Supplier D
W32A 500 560 530 550
Y33A 109 121 115 119
W32B 27 33 30 32
X33B 230 248 239 245
Z32A 324 372 348 364

Our objective is to obtain our required level of parts inventory at the minimum cost. Thus our objective function is-

Min((Xij*Cij)

The constraints are-

From number of parts needed-

X1A + X1B + X1C + X1D = 6000

X2A+X2B+X2C+X2D = 32000

X3A+X3B+X3C+X3D=22000

X4A+X4B+X4C+X4D=6000

X5A+X5B+X5C+X5D= 27000

From the maximum capability of each supplier-

X1A+X2A+X3A+X4A+X5A <= 26400

X1B+X2B+X3B+X4B+X5B <= 25600

X1C+X2C+X3C+X4C+X5C <= 3000

X1D+X2D+X3D+X4D+X5D <= 41000

As the number of parts can't be negative, Xij >= 0

The solver screenshots are-

Considering the levels of service of the suppliers, what is the minimum total procurement cost for GetFit? = 25407800 (from Objective cell J1)

Please Like & Provide your reviews in comments. :-)